OpenStudy (anonymous):

Solve for x: log2(x−4)log2(x−3)=1

OpenStudy (anonymous):

\[\log_2(x−4)\log_2(x−3)=1\]Is this the correct problem?

OpenStudy (anonymous):

ah, I got it. I forgot this law. Thank you! :)

OpenStudy (anonymous):

I don't see how that solves the question. (x -4)(x - 3) = 2 would give x = 2 or 5. Substitute 2 or 5 into the original equation doesn't give 1. Am I doing something wrong here? Anyway how can you raise each term like this? \[ a * b = c \rightarrow 2^{a} * 2^{b} = 2^{c}\]since \[2^{a} * 2^{b} = 2^{c} \rightarrow 2^{a + b} = 2^{c} \rightarrow a+b = c\]